OpenStudy (anonymous):

Find the product of (x + 7)2. x2 - 14x + 49 x2 + 14x + 49 x2 - 49 x2 + 49

OpenStudy (gorv):

just open the square or (x+7)^2=(x+7)(x+7) now expand =x*(x+7)+7*(x+7) =x^2+7x+7x+49 =x^2+14x+49

OpenStudy (anonymous):

ok thanks :D

OpenStudy (anonymous):

could you help me with a few more?

OpenStudy (gorv):

post them seprately i will help

OpenStudy (anonymous):

ok